# Orbifold modifications of complex analytic varieties

### v4

@@ 8. Proof of the modification theorem
We can apply Lemma 8.1 on each $X_{e}$, and obtain a projective bimeromorphic morphism $\varphi_{e}\colon Y_{e}\to Z_{e}$ as in the Lemma 8.1. The indeterminacy locus of $(\varphi_{e})^{-1}$ is contained in $h^{-1}(S_{e})$, hence is compact. Since there is a $\varphi_{e}$-ample and $\varphi_{e}$-exceptional divisor, by applying Lemma 2.3 for several times, we deduce that there is a projective bimeromorphic morphism $\varphi\colon Y\to Z$ such that its restriction over $Z_{e}$ is equal to $\varphi_{e}$, and that it is an isomorphism elsewhere. In particular, $Y$ has quotient singularities. Let $f\colon Y\to X$ be natural morphism. Then $f$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of $S_{1},...,S_{r}$. This completes the proof of the theorem. ∎
9. The Bogomolov-Gieseker inequality
In this section, we will prove Theorem 1.3. The notion of Kähler spaces was introduced in [Gra62]. For Kähler orbifolds and orbifold coherent sheaves, see for example [Fau22, Section 2], [Wu23, Section 2] or [DO23, Section 3.1].
If $X$ is a compact complex analytic variety of dimension $n$, then there is a natural isomorphism $H_{2n}(X,\mathbb{R})\cong H_{2n}^{BM}(X,\mathbb{R})$, where $H_{2n}$ is the singular homology and $H_{2n}^{BM}$ is the Borel-Moore homology. We note that $H_{2n}^{BM}(X,\mathbb{R})\cong\mathbb{R}$ and it is generated by a canonical element $[X]$, called the fundamental class of $X$. For more details on Borel-Moore homology, we refer to [Ful98, Chapter 19] and the references therein. As a consequence, for cohomology classes $\sigma_{1},...,\sigma_{k}\in H^{\bullet}(X,\mathbb{R})$, such that the sum of their degrees is equal to $2n$, we can define the following intersection number,
$\sigma_{1}\cdots\sigma_{k}:=(\sigma_{1}\smallsmile\cdots\smallsmile\sigma_{k})\smallfrown[X]\in\mathbb{R}.$
When $X$ has only quotient singularities, this intersection form can be identified with the orbifold Poincaré duality, see [Sat56, Theorem 3].
9.A. Orbifold Chern classes
We will explain the orbifold Chern classes $\hat{c}_{k}$ for $k=1,2$ in the statement of Theorem 1.3. We also refer to [GK20, Section 5] for a similar discussion.
Let $X$ be a compact complex analytic variety of dimension $n$, which has quotient singularities in codimension 2. Let $\mathcal{F}$ be a coherent reflexive sheaf on $X$. Then there is a Zariski open subset $X^{\circ}$ of $X$, whose complement has codimension at least 3, such that there is a standard complex orbifold $\mathfrak{X}=(X_{j},G^{\prime}_{j})$ whose quotient space is $X^{\circ}$. Being standard means that the action of $G^{\prime}_{j}$ is free in codimension 1. Furthermore, since reflexive sheaves on complex manifolds are locally free in codimension 2, up to shrinking $X^{\circ}$, we can assume that $\mathcal{F}$ induces an orbifold vector bundle $\mathcal{F}_{\mathrm{orb}}$ on $\mathfrak{X}$, see [DO23, Remark 3.4.(5)].
Let $f\colon Y\to X$ be an orbifold modification as in Theorem 1.2. Then $Y$ has quotient singularities. Let $\mathfrak{Y}=(Y_{i},G_{i})$ be the standard orbifold structure on $Y$, and let $\mathcal{E}=(f^{*}\mathcal{F})/(\mathrm{torsion})$. Then $\mathcal{E}$ induces a reflexive coherent orbifold sheaf $\mathcal{E}_{\mathrm{orb}}$ on $\mathfrak{Y}$, see for example [DO23, Remark 3.4.(5)]. By shrinking $X^{\circ}$, we may assume that $f$ is an isomorphism over it.
By [DO23, Theorem 3.10], there is a functorial projective bimeromorphic morphism $p_{i}\colon Z_{i}\to Y_{i}$, such that $Z_{i}$ is smooth, that $\mathcal{H}_{i}:=p_{i}^{*}\mathcal{E}_{i}/(\mathrm{torsion})$ is locally free, and that the indeterminacy locus of $p_{i}^{-1}$ has codimension at least 3. The functoriality implies that there is a canonical action of $G_{i}$ on $Z_{i}$ such that $p_{i}$ is $G_{i}$-equivariant. Furthermore, the $(Z_{i},G_{i})$’s induce a complex orbifold $\mathfrak{Z}$ with quotient space $Z$. The collection $(\mathcal{H}_{i})$ defines an orbifold vector bundle $\mathcal{H}_{\mathrm{orb}}$ on $\mathfrak{Z}$. Therefore, we have well-defined orbifold Chern classes $\hat{c}_{1}(\mathcal{H}_{\mathrm{orb}})$ and $\hat{c}_{2}(\mathcal{H}_{\mathrm{orb}})$ in $H^{\bullet}(Z,\mathbb{R})$, see [Bla96, Section 2] or [DO23, Definition 3.3]. Let $q\colon Z\to X$ be the natural morphism. Then the indeterminacy locus of $q^{-1}$ has codimension at least 3. We can now define the orbifold Chern classes $\hat{c}_{1}(\mathcal{F})$, $\hat{c}_{2}(\mathcal{F})$ and $\hat{c}_{1}(\mathcal{F})^{2}$.
Definition 9.1.
The orbifold Chern class $\hat{c}_{1}(\mathcal{F})$, $\hat{c}_{2}(\mathcal{F})$ and $\hat{c}_{1}(\mathcal{F})^{2}$ are defined as linear forms on $H^{\bullet}(X,\mathbb{R})$, so that for any class $\sigma\in H^{\bullet}(X,\mathbb{R})$, we have
$\hat{c}_{k}(\mathcal{F})\cdot\sigma=\hat{c}_{k}(\mathcal{H}_{\mathrm{orb}})\cdot q^{*}\sigma$ and $\hat{c}_{1}(\mathcal{F})^{2}\cdot\sigma=\hat{c}_{1}(\mathcal{H}_{\mathrm{orb}})^{2}\cdot q^{*}\sigma,$
where $k=1,2$.
To see that these classes are well-defined, it is equivalent to show that the intersection numbers are independent of the choice of the modification $q\colon Z\to X$. We first observe the following fact.
@@ 9.A. Orbifold Chern classes
Let $X$ be compact complex analytic variety of dimension $n$ and let $\rho\colon\widetilde{X}\to X$ be a proper bimeromorphic morphism. Let $\alpha_{1},\alpha_{2}\in H^{2k}(\widetilde{X},\mathbb{R})$. Assume that there is a closed analytic subset $V\subseteq X$ of codimension at least $k+1$, such that $\alpha_{1}|_{\widetilde{X}^{\circ}}=\alpha_{2}|_{\widetilde{X}^{\circ}}$, where ${\widetilde{X}^{\circ}}=\widetilde{X}\setminus\rho^{-1}(V)$. Then for any class $\sigma\in H^{2n-2k}(X,\mathbb{R})$, we have
$\alpha_{1}\cdot\rho^{*}\sigma=\alpha_{2}\cdot\rho^{*}\sigma.$
@@
Let $E=\rho^{-1}(V)$. Then there is some class $\delta\in H^{2k}(\widetilde{X},\widetilde{X}\setminus E,\mathbb{R})$ such that $\alpha_{1}-\alpha_{2}$ is equal to the image of $\delta$ in $H^{2k}(\widetilde{X},\mathbb{R})$. It follows that
$(\alpha_{1}-\alpha_{2})\cdot\rho^{*}\sigma$ $=$ $\delta\cdot\rho^{*}\sigma$
$=$ $(\rho^{*}\sigma\smallsmile\delta)\smallfrown[\widetilde{X}]$
$=$ $\rho^{*}\sigma\smallfrown(\delta\smallfrown[\widetilde{X}]).$
We note that $\delta\smallfrown[\widetilde{X}]\in H_{2n-2k}(E,\mathbb{R})$. Then $\rho_{*}(\delta\smallfrown[\widetilde{X}])\in H_{2n-2k}(V,\mathbb{R})$. By assumption, we have $\operatorname{dim}V\leq n-k-1$. Hence $\rho_{*}(\delta\smallfrown[\widetilde{X}])=0$. By the projection formula, we have
$\rho^{*}\sigma\smallfrown(\delta\smallfrown[\widetilde{X}])=\sigma\smallfrown\rho_{*}(\delta\smallfrown[\widetilde{X}])=0.$
This completes the proof of the lemma. ∎
We are in position to prove the following statement.
Proposition 9.3.
The intersection numbers in Definition 9.1 are independent of the choice of $Z$.
Proof.
We assume the following setting. Let $q_{j}\colon Z_{j}\to X_{j}$ be proper bimeromorphic morphisms with $j=1,2$. There is a closed analytic subset $V\subseteq X$ of codimension at least 3, such that $X^{\circ}:=X\setminus V$ has quotient singularities. The restriction of the reflexive coherent sheaf $\mathcal{F}$ on $X^{\circ}$ induces an orbifold vector bundle $\mathcal{F}_{\mathrm{orb}}$ on the standard orbifold structure of $X^{\circ}$. Let $Z_{j}^{\circ}=q_{j}^{-1}(X^{\circ})$. Then $q_{j}$ is an isomorphism on $Z_{j}^{\circ}$. There is an orbifold $\mathfrak{Z}_{j}$ whose quotient space is equal to $Z_{j}$, such that $\mathfrak{Z}_{j}$ is standard over $Z_{j}^{\circ}$. There is an orbifold vector bundle $\mathcal{H}_{\mathrm{orb},j}$ on $\mathfrak{Z}_{j}$, whose restriction over $Z_{j}^{\circ}$ is isomorphic to $\mathcal{F}_{\mathrm{orb}}$ over $X^{\circ}$. Let $\alpha_{j}\in\{\hat{c}_{1}(\mathcal{H}_{\mathrm{orb},j}),\hat{c}_{2}(\mathcal{H}_{\mathrm{orb},j}),\hat{c}_{1}(\mathcal{H}_{\mathrm{orb},j})^{2}\}$, such that $\alpha_{1}$ and $\alpha_{2}$ are the same type of characteristic class.
To prove the proposition, it is enough to show that, for any $\sigma\in H^{\bullet}(X,\mathbb{R})$, the following equality holds
$\alpha_{1}\cdot q_{1}^{*}\sigma=\alpha_{2}\cdot q_{2}^{*}\sigma.$
We note that there is a natural bimeromorphic map $\varphi\colon Z_{1}\dashrightarrow Z_{2}$ over $X$. Let $\widetilde{X}\subseteq Z_{1}\times Z_{2}$ be the closure of the graph of $\varphi$, and let $p_{1}\colon\widetilde{X}\to Z_{1}$, $p_{2}\colon\widetilde{X}\to Z_{2}$ and $\rho\colon\widetilde{X}\to X$ be the natural morphisms. Then we have
$\alpha_{j}\cdot q_{j}^{*}\sigma=p_{j}^{*}\alpha_{j}\cdot\rho^{*}\sigma$
for $j=1,2$. We notice that $p_{1}$, $p_{2}$ and $\rho$ are isomorphisms over $X^{\circ}$. Hence if $\widetilde{X}^{\circ}=\rho^{-1}(X^{\circ})$, then we have
$(p_{1}^{*}\alpha_{1})|_{\widetilde{X}^{\circ}}=(p_{2}^{*}\alpha_{2})|_{\widetilde{X}^{\circ}}.$
By applying Lemma 9.2, we obtain that $\alpha_{1}\cdot q_{1}^{*}\sigma=\alpha_{2}\cdot q_{2}^{*}\sigma.$ This completes the proof of the proposition. ∎
Remark 9.4.
We gather some properties on this notion of Chern classes, which can be derived directly from the definition and Lemma 9.2.
  1. (1)
    If $X$ has quotient singularities and if $\mathcal{F}$ induces an orbifold vector bundle $\mathcal{F}_{\mathrm{orb}}$ on the standard orbifold structure over $X$, then the Chern classes of $\mathcal{F}$ in Definition 9.1 coincides with the linear forms on $H^{\bullet}(X,\mathbb{R})$ induced by the orbifold Chern classes of $\mathcal{F}_{\mathrm{orb}}$.
  2. (2)
    Let $X^{\circ}\subseteq X$ be an open subset with quotient singularities, such that $\mathcal{F}|_{X^{\circ}}$ induces an orbifold vector bundle $\mathcal{F}_{\mathrm{orb}}$ on the standard orbifold structure on $X^{\circ}$, and that $X\setminus X^{\circ}$ has codimension at least 3. Let $\Delta\in\{\hat{c}_{1},\hat{c}_{2},\hat{c}_{1}^{2}\}$. If there is some class $\beta\in H^{\bullet}(X,\mathbb{R})$, whose restriction on $X^{\circ}$ is equal to $\Delta(\mathcal{F}_{\mathrm{orb}})$, then $\Delta(\mathcal{F})$ is the same as the linear form on $H^{\bullet}(X,\mathbb{R})$ defined by the intersection with $\beta$.
  3. (3)
    Let $M=H_{1}\cap\cdots\cap H_{n-2}$ be the complete intersection surface of basepoint-free Cartier divisors in general position. Then $M$ has quotient singularities, $\mathcal{F}|_{M}$ is reflexive and inducing an orbifold vector bundle on the standard orbifold structure of $M$, and we have
    $\Delta(\mathcal{F})\cdot c_{1}(H_{1})\cdots c_{1}(H_{n-2})=\Delta(\mathcal{F}|_{M}),$
    where $\Delta$ is either $\hat{c}_{2}$ or $\hat{c}_{1}^{2}$.
With the first Chern class in Definition 9.1, we can extend the notion of slope stability as follows. Let $X$ be a compact complex analytic variety of dimension $n$, which has quotient singularities in codimension 2. Let $\mathcal{F}$ be a coherent reflexive sheaf on $X$. For a cohomology class $\alpha\in H^{2}(X,\mathbb{R})$, we can define the slope
$\mu_{\alpha}(\mathcal{F})=\frac{\hat{c}_{1}(\mathcal{F})\cdot\alpha^{n-1}}{\mathrm{rank}\,F}.$
Then $\mathcal{F}$ is called $\alpha$-semistable if for any nonzero coherent subsheaf $\mathcal{E}\subseteq\mathcal{F}$, we have $\mu_{\alpha}(\mathcal{E})\leq\mu_{\alpha}(\mathcal{F})$. It is called $\alpha$-stable if the inequality is strict whenever $\mathcal{E}$ has smaller rank. From the item (2) of Remark 9.4, we see that this definition coincides with the classic one, if $X$ satisfies that every reflexive coherent sheaf of rank 1 on it has a positive reflexive power which is locally free.
9.B. Proof of the Bogomolov-Gieseker inequality
We complete the proof of Theorem 1.3 in this subsection.
Proof of Theorem 1.3.
Let $f\colon Y\to X$ be an orbifold modification as in Theorem 1.2. Then $Y$ has quotient singularities. It follows that $Y$ is the quotient space of an orbifold $\mathfrak{Y}=(Y_{i},G_{i})_{i\in I}$, such that the action of $G_{i}$ is free in codimension 1. Since $f$ is projective, $Y$ is a compact Kähler variety. Let $\mathcal{E}=(f^{*}\mathcal{F})/(\mathrm{torsion})$. Then $\mathcal{E}$ induces a reflexive coherent orbifold sheaf $\mathcal{E}_{\mathrm{orb}}$ on $\mathfrak{Y}$, see for example [DO23, Remark 3.4.(5)]. We note that $\mathcal{E}$ is stable with respect to $f^{*}\omega$. By [DO23, Remark 3.4.(5)] again, we deduce that $\mathcal{E}_{\mathrm{orb}}$ is stable with respect to the class of $f^{*}\omega$.
By [DO23, Theorem 3.10], there is a functorial projective bimeromorphic morphism $p_{i}\colon Z_{i}\to Y_{i}$, such that $Z_{i}$ is smooth, that $\mathcal{H}_{i}:=p_{i}^{*}\mathcal{E}_{i}/(\mathrm{torsion})$ is locally free, and that the indeterminacy locus of $p_{i}^{-1}$ has codimension at least 3. In particular, there is a canonical action of $G_{i}$ on $Z_{i}$ such that $p_{i}$ is $G_{i}$-equivariant, and that the $(Z_{i},G_{i})$’s induce a complex orbifold $\mathfrak{Z}$ with quotient space $Z$. The collection $(\mathcal{H}_{i})$ defines an orbifold vector bundle $\mathcal{H}_{\mathrm{orb}}$ on $\mathfrak{Z}$. There is an induced projective bimeromorphic morphism $p\colon Z\to Y$. In particular, $Z$ is a Kähler variety. Then there is an orbifold Kähler form $\eta$ on $\mathfrak{Z}$, see [Wu23, Lemma 1]. Let $q\colon Z\to X$ be the natural morphism. We note that $q^{*}\omega$ can be viewed as a semipositive orbifold $(1,1)$-form on $\mathfrak{Z}$. Moreover, $\mathcal{H}_{\mathrm{orb}}$ is stable with respect to $q^{*}\omega$ for $\mathcal{E}_{\mathrm{orb}}$ is stable with respect to $f^{*}\omega$. By the same argument of [DO23, Lemma 3.16], we deduce that, for any $\varepsilon>0$ small enough, $\mathcal{H}_{\mathrm{orb}}$ is stable with respect to the following orbifold Kähler form on $\mathfrak{Z}$
$\alpha_{\varepsilon}:=q^{*}\omega+\varepsilon\eta.$
By [Fau22, Theorem 1], $\mathcal{H}_{\mathrm{orb}}$ admits an orbifold Hermitian-Einstein metric with respect to $\alpha_{\varepsilon}$. Then, as in [Kob14, Theorem 4.4.7], we have
$\Big{(}2r\hat{c}_{2}(\mathcal{H}_{\mathrm{orb}})-(r-1)\hat{c}_{1}(\mathcal{H}_{\mathrm{orb}})\Big{)}\cdot[\alpha_{\varepsilon}]^{n-2}\geq 0.$
The limit of the LHS above, when $\varepsilon$ tends to 0, is equal to
$\Big{(}2r\hat{c}_{2}(\mathcal{H}_{\mathrm{orb}})-(r-1)\hat{c}_{1}(\mathcal{H}_{\mathrm{orb}})\Big{)}\cdot[q^{*}\omega]^{n-2}$
$=$ $\Big{(}2r\hat{c}_{2}(\mathcal{F})-(r-1)\hat{c}_{1}(\mathcal{F})^{2}\Big{)}\cdot[\omega]^{n-2}$
This completes the proof of the theorem. ∎

### v6

@@ 8. Proof of the modification theorem
We can apply Lemma 8.1 on each $X_{e}$, and obtain a projective bimeromorphic morphism $\varphi_{e}\colon Y_{e}\rightarrow Z_{e}$ as in the Lemma 8.1. The indeterminacy locus of $(\varphi_{e})^{-1}$ is contained in $h^{-1}(S_{e})$, hence is compact. Since there is a $\varphi_{e}$-ample and $\varphi_{e}$-exceptional divisor, by applying Lemma 2.3 for several times, we deduce that there is a projective bimeromorphic morphism $\varphi\colon Y\rightarrow Z$ such that its restriction over $Z_{e}$ is equal to $\varphi_{e}$, and that it is an isomorphism elsewhere. In particular, $Y$ has quotient singularities. Let $f\colon Y\rightarrow X$ be natural morphism. Then $f$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of $S_{1},...,S_{r}$. This completes the proof of the theorem. ∎
Remark 8.2.
We may improve the construction above so that there is some $f$-exceptional and $f$-ample divisor, by the method of [KW24]. Let $f_{1}\colon Y_{1}\rightarrow X$ be a morphism constructed in Theorem 1.2, and let $\mathcal{H}$ be a $f_{1}$-ample line bundle. We set $\mathcal{L}=((f_{1})_{*}\mathcal{H})^{*}$ and $Y_{2}=\mathrm{Proj}\bigoplus_{m\geq 0}\mathcal{L}^{\otimes m}$. Let $W$ be the normalization of the main component of $Y_{1}\times_{X}Y_{2}$. Then the indeterminacy locus of the natural map $X\dashrightarrow W$ is contained to the one of $(f_{1})^{-1}$. We recall that $S_{1},...,S_{r}$ are the irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$ which have codimension $2$. We apply Lemma 2.18 to $W$ with respect to the strict transforms of $S_{1},...,S_{r}$, and obtain a morphism $V\rightarrow W$. Afterwards, we apply Theorem 1.2 to $V$ to obtain a modification $Y\rightarrow V$. Then we claim that the natural morphism $f\colon Y\rightarrow X$ satisfies the properties of Theorem 1.2. In fact, assume by contradiction that there is a component $Z$ of the indeterminacy locus of $f^{-1}$, which has codimension $2$. Then the strict transform $Z_{W}$ of $Z$ in $W$ is well-defined. Since the indeterminacy locus of $V\dashrightarrow Y$ has codimension at least 3, we see that $Z_{W}$ is contained in the indeterminacy locus of $W\dashrightarrow V$. By the construction of $V$, see Lemma 2.18 (2), $Z$ is not equal to anyone of the subvarieties $S_{1},...,S_{r}$. Therefore, $Z$ meets the smooth locus of $X$, and hence $Z_{W}$ meets $W_{\operatorname{\textsubscript{\rm sm}}}$. This contradicts Lemma 2.18 (1). In the end, by using Lemma 8.3 below, we can adapt the argument of [KW24, Lemma 8] to show that, there is some $f$-exceptional and $f$-ample divisor.
@@
Let $f\colon Y\rightarrow X$ be a projective bimeromorphic morphism between compact normal complex analytic singularities. Assume that $\mathcal{L}$ is a line bundle on $Y$, such that $\mathcal{L}|_{Y\setminus\operatorname{Ex}(f)}$ is trivial, where $\operatorname{Ex}(f)$ is the $f$-exceptional locus. Then there is some $f$-exceptional Cartier divisor $D$ such that $\mathcal{L}\cong\mathcal{O}_{Y}(D)$.
@@
We set $\mathcal{I}=f_{*}\mathcal{L}$. Then $\mathcal{I}$ is a torsion-free coherent sheaf on $X$ and $\mathcal{I}^{**}\cong\mathcal{O}_{X}$. Hence we may assume that $\mathcal{I}$ is an ideal sheaf on $X$, whose cosupport is contained in $f(\operatorname{Ex}(f))$. There is a natural morphism $f^{*}\mathcal{I}\rightarrow\mathcal{L}$, which factors through $\mathcal{J}:=f^{*}\mathcal{I}/(\mathrm{torsion})$. We note that $\mathcal{J}$ is indeed equal to the ideal sheaf $f^{-1}\mathcal{I}\cdot\mathcal{O}_{Y}$, whose cosupport is contained in $\operatorname{Ex}(f)$.
We first assume that there is some effective Cartier divisor $\Gamma$, whose support is equal to $\operatorname{Ex}(f)$. Then there is some integer $j>0$, such that $\mathcal{O}_{Y}(-j\Gamma)$ is contained in $\mathcal{J}$. We hence obtain an injective morphism $\mathcal{O}_{Y}\rightarrow\mathcal{L}\otimes\mathcal{O}_{Y}(j\Gamma)$. By taking its dual, we obtain an injective morphism $\mathcal{L}^{*}\otimes\mathcal{O}_{Y}(-j\Gamma)\rightarrow\mathcal{O}_{Y}$. It follows that $\mathcal{L}^{*}\otimes\mathcal{O}_{Y}(-j\Gamma)\cong\mathcal{O}_{Y}(-E)$ for some effective $f$-exceptional divisor $E$ supported inside $\operatorname{Ex}(f)$. Thus $\mathcal{L}\cong\mathcal{O}_{Y}(E-j\Gamma)$.
In general, let $g\colon Z\rightarrow Y$ be a desingularization, let $h=f\circ g$ and let $\mathcal{M}=g^{*}\mathcal{L}$. We may assume that $\operatorname{Ex}(h)$ is pure of codimension 1. From the previous paragraph, there is some $h$-exceptional divisor $D_{1}$ on $Z$, such that $\mathcal{M}\cong\mathcal{O}_{Z}(D_{1})$. Let $D=g_{*}D_{1}$. Then $D$ is supported in $\operatorname{Ex}(f)$, and the reflexive sheaf $\mathcal{O}_{Y}(D)\otimes\mathcal{L}^{*}$ is isomorphic to $\mathcal{O}_{Y}$ on $Y\setminus g(\operatorname{Ex}(g))$. This implies that $\mathcal{L}\cong\mathcal{O}_{Y}(D)$, and completes the proof of the lemma. ∎